OpenStudy (anonymous):

whats the 43 term? 13.1,3.1,-6.9,-16.9

OpenStudy (amistre64):

13.1 -3.1 ----- 10 3.1 6.9 ---- 10 its adding 10s

OpenStudy (amistre64):

well. subtracting 10 :)

OpenStudy (amistre64):

A{n} = 23.1-10n

OpenStudy (anonymous):

amistre u r making a mistake !!! it should be 3.1 - 13.1 !!!!!! to ding the common difference

OpenStudy (amistre64):

A{43} = 23.1 - 10(43) A{43} = 23.1- 430 430.0 - 23.1 ------ 406.9 - hmm.... you might have a point there A{1} = 23.1 - 10(1) = 13.1 A{2} = 23.1 - 10(2) = 3.1

OpenStudy (anonymous):

sorry correct ans is -406.9

OpenStudy (amistre64):

well, if im wrong; we are wrong together :) cause I got -406.9 too ;)

OpenStudy (anonymous):

No it is 100% correct. the first time around i pressed x instead of + !!!

OpenStudy (anonymous):

the choices is a. 35.1, b.35.5 c.36.3

OpenStudy (amistre64):

then your choices are not to this question ... check the question and see if there is a typo in it

OpenStudy (anonymous):

yeah i feel the same if the choices r as u wrote....

OpenStudy (anonymous):

what about find the 43rd term 19.5, 19.9, 0.3, 20.7, a.35.1 b.35.5 c.35.9 d.36.3